# Conceptualization of this study, Methodology, Software

### v1

@@ Abstract
Hirschsprung Disease is characterized by the absence of ganglion cells in the myenteric plexus. Therefore, their correct identification is crucial for diagnosing Hirschsprung disease. We introduce a three-stage segmentation framework based on a Vision Transformer (ViT-B/16) that mimics the pathologist’s diagnostic approach. The framework sequentially segments the muscularis propria, delineates the myenteric plexus, and identifies ganglion cells within anatomically valid regions. 30 whole-slide images of colon tissue were used, each containing expert manual annotations of muscularis, plexus, and ganglion cells at varying levels of certainty. A 5-fold cross-validation scheme was applied to each stage, along with resolution-specific tiling strategies and tailored postprocessing to ensure anatomical consistency. The proposed method achieved a Dice coefficient of 89.9% and a Plexus Inclusion Rate of 100% for muscularis segmentation. Plexus segmentation reached a recall of 94.8%, a precision of 84.2% and a Ganglia Inclusion Rate of 99.7%. For high-certainty ganglion cells, the model achieved 62.1% precision and 89.1% recall, while joint certainty scores yielded 67.0% precision. These results indicate that ViT-based models are effective at leveraging global tissue context and capturing cellular morphology at small scales, even within complex histological tissue structures. This multi-stage methodology has great potential to support digital pathology workflows by reducing inter-observer variability and assisting in the evaluation of Hirschsprung disease. The clinical impact will be evaluated in future work with larger multi-center datasets and additional expert annotations.
@@
Hirschsprung disease (HD), a congenital birth defect disorder defined by the absence of ganglionic cells within the myenteric plexus of the intestinal tract [7], rapidly manifesting after birth, affects 1/5000 neonatal patients worldwide [1]. Lack of ganglionic cells, caused by defective migration, proliferation, and differentiation of neural crest cells during developmental periods throughout gestation. Thus, leading to obstruction of bowel movements caused by malformation of colon nerve cells and therefore leading to symptoms of HD, including severe constipation and signs of intestinal obstruction [2]. Thus, accurate identification of ganglionic cells in the colon is vital for confirmation of HD.
@@
After the onset of symptoms, traditional methods for diagnosing HD include a contrast enema radiograph, in which a barium contrast is inserted into the child’s rectum and x-ray images of the colon are taken to visualize the structure of different colon segments, thereby confirming the presence of malformed intestines. A second popular method for diagnosing HD is a rectal biopsy, in which a section of tissue is extracted from the rectum and examined on a Whole Slide Image (WSI) by a trained pathologist. This method involves quantitatively assessing the presence of ganglionic cells. Although both are typically considered the gold standard for confirming HD, they have inherent limitations, such as sampling error, inter- and intra-rater variability among pathologists [8, 10], and excessive time consumption. Highlighting the importance of automating the identification of ganglionic cells to confirm HD diagnosis.
@@
The current advances in computational imaging and WSI have enabled Artificial Intelligence (AI) methods to automate the classification, detection, and segmentation of medical image data across various organ systems. For instance, several Deep Learning (DL) algorithms have been utilized for computer-aided diagnosis, attempting to classify breast cancer lesions [9] using WSIs [3] and human aortic and cardiac regions [5]. Other algorithms [4] have been deployed to segment images of bronchoscopes, further identifying lung-related diseases. Furthermore, object segmentation has also been widely applied to various medical images and diseases. Specifically, classifying the potential onset of autism spectrum disorder (ASD), including additional indicators of ASD and irregular brain structure [24, 25, 30]. Leveraging DL computer vision algorithms for identifying the presence of ganglion cells does seem promising. However, segmentation is required because ganglion cells are dispersed across the myenteric plexus layer, which lies within the muscularis layer of the colon. Segmenting the different colon layers enables later identification of ganglion cells, or their absence, thereby further confirming the presence of HD.
@@
Recent attempts [11, 12, 13, 14, 15] have been made to segment the muscularis layer, which encloses the myenteric plexus, first utilizing a shallow machine learning model, K-means clustering [14]. This was then followed by deploying a Convolutional Neural Network (CNN), which produced a far greater accuracy and plexus inclusion rate when segmenting the muscularis layer [13, 14]. In [14], a colour-thresholding algorithm was then utilized for segmenting the myenteric plexus region and effectively detecting Carletinin-positive ganglia inside plexus regions. Finally, ganglia were segmented from each plexus region using intensity thresholding on the previously segmented Calretinin signal. This was followed by deploying a Linear Discriminant Analysis (LDA) model to identify true-ganglia from false-ganglia [14]. Overall, the results suggest strong discriminative ability, offering a promising avenue for automated DL. Furthermore, other attempts have been made [6] that leverage the U-Net, a CNN algorithm, to similarly detect the presence of ganglion cells in the myenteric plexus, using WSIs from biopsy tissues. M. Duci et. al., in [6], demonstrated superior innate capabilities of the CNN architecture, especially given its ability to extract low-level and high-level features throughout different layers, in addition to capturing local relationships through convolutions. Overall, the algorithm is able to learn rich, essential semantic features from the colon layers in the WSI dataset. Although U-Net is appropriate for learning pixel-wise local representations, it cannot learn long-range, global contextual relationships across different regions of the image, which could be critical for detecting ganglion cells dispersed throughout the WSI.
@@
Given its success in Natural Language Processing (NLP) tasks, Vision Transformer (ViT) [17] have achieved superior performance to CNNs in computer vision-related tasks, when utilized on general image datasets in addition to medical image data [5], as shown in the myenteric plexus classification study in [16]. This is possible because ViTs have inherent self-attention, which enables them to model long-range dependencies and capture global contextual relationships across all parts of an image. Thereby, recognizing patterns that span large areas, such as tissue structure in WSI histopathology.
In this study, we propose a hierarchical segmentation pipeline utilizing ViTs to segment the muscularis layer, which encloses the myenteric plexus. This is followed by segmenting the myenteric plexus layer and finally identifying the presence of ganglion cells, typically found in the myenteric plexus, to further confirm the onset of HD. The study used 30 WSIs of colon sections, manually annotated by a pediatric pathologist. The primary metrics used for evaluating segmentation performance of the muscularis include Dice-Sørensen coefficient (Dice coefficient), precision, recall, and Plexus Inclusion Rate (PIR), ultimately assessing the similarity between the ground truth dataset and the prediction produced. For evaluating the myenteric plexus, evaluation metrics include precision, recall, and Ganglia Inclusion Rate (GIR). Finally, considering the presence of ganglion cells, it was also necessary to assess recall and precision.
@@ 2.1 Dataset
The dataset used in this research study originates from the Children’s Hospital of Eastern Ontario (CHEO), comprising 30 WSIs corresponding to colon sections from 26 different patients, all of whom were diagnosed with HD. All high-resolution images were extracted from prepared tissue slides after being scanned using the Aperio ScanScope CS (Aperio Technologies) at a 20 magnification level (0.50 /pixel). All 30 WSIs were saved in an SVS format. For each layer, there exists ground truth annotations that a neonatal pathologist manually segmented (Figure 1). For instance, the muscularis propria was manually delineated to represent the ground-truth segmentation, the myenteric plexus region was roughly manually segmented (a visually noticeable amount of tissue around the plexus regions was also included), followed by ganglion cells, which were also roughly segmented. A confidence level accompanies each ganglion cell annotation. A high-confidence level indicates strong certainty that the annotated object is a ganglion cell (Figure 2C). In contrast, a low-confidence level reflects that the object is believed to be a ganglion cell but with some uncertainty (Figure 2D), due to the difficulty in recognizing and selecting physiological features of a ganglion cell [15]. An overview of these tissue regions at different magnification levels is shown in Figure 2.
@@ 2.2.1 Muscularis Propria
Preprocessing for the muscularis propria region involved three steps. First, WSIs were down-sampled by a factor of 10 from the original 20 (0.50 /pixel) magnification, to the final resolution of 5 /pixel. The lower magnification was chosen because the muscularis propria occupies a large, continuous region of the colon, so a more zoomed-out view of the tissue was better able to capture its global structure. Second, colour normalization was applied using the Macenko normalization method [18] to correct for staining variation. Finally, the 30 WSIs were tiled into 224224-pixel patches (1000 tiles per WSI, for a total of 30,000 tiles), enabling the model to learn the characteristic texture, morphology, and appearance of the muscularis layer across the slide. An example of the muscularis ground truth annotations used in this stage is presented in Figure 1B.
@@ 2.2.2 Myenteric Plexus
Preprocessing for the myenteric plexus region followed the same general procedure but was adjusted based on the plexus’s anatomy. The WSIs were down-sampled by a factor of 4 from 20 (0.50 /pixel) to a final magnification of 5 (2.50 /pixel). This level was chosen as it offers a compromise between resolution and spatial context, preserving finer details while still providing a sufficiently large field of view. Macenko normalization [18] was applied to correct for staining variation. The WSIs were tiled into 224224-pixel patches, and the 30,000 tiles were selected if they contained at least one pixel within the muscularis propria region. This step was included because the myenteric plexus is strictly contained within the muscularis layer, so there is no need to tile other tissue regions. A reference example of plexus ground truth annotations is shown in Figure 1C, which shows how the plexus ground truth is exclusively within the muscularis segmentation, and in 3B.
@@ 2.2.3 Ganglion Cells
Preprocessing for ganglion cell segmentation followed the same general workflow but was modified to be suitable for cellular-level analysis. The WSIs were down-sampled by a factor of 2 from 20 (0.50 /pixel) to a final magnification of 10 (1.0 /pixel). The higher resolution provides a closer view of the tissue, preserving the fine morphological features required to identify ganglion cells. Macenko normalization [18] was used to correct for staining variation. The images were tiled into 224224 pixel patches, and only tiles containing at least one pixel from the myenteric plexus region were selected. Since ganglion cells are exclusively found within the plexus, this ensured that the model was trained only on relevant regions.
@@ 2.3.1 Model Architecture
The segmentation framework was based on a ViT-B/16 backbone [17]. Specifically, each 224224 tile was divided into fixed-sized patches that were linearly projected to patch embeddings and then fed to the transformer encoder. The encoder generated a global contextual representation shared among the three downstream segmentation tasks. To enable the hierarchical workflow, three task-specific decoders were employed to independently generate prediction maps for the muscularis, plexus, and ganglion. For a given task, the decoder received the encoded representations and produced a dense pixel output for the target layer. An overview of the complete three-stage pipeline is provided in Figure 4.
@@ 2.3.2 Muscularis Propria
To ensure robust segmentation of the muscularis propria and reduce the risk of overfitting, especially given the small dataset size, a 5-fold cross-validation strategy was employed. During training, the dataset was split into five equal folds, each comprising 6 WSIs (6000 tiles) for validation. This process was applied iteratively, so that each WSI was used both as a training set in some folds and as a test set in others. Thereby, making better use of the dataset. To increase dataset diversity, both sets of tiles were augmented using four methods: scaling, vertical flipping, horizontal flipping, and random rotation. The scaling operation simulated zoom effects by randomly enlarging or shrinking the visible region before resizing the tile back to 224224 pixels. This augmentation enabled the model to learn from tiles representing different effective magnifications. Rotation entailed rotating the tiles by 90; horizontal and vertical flipping consisted of flipping the images by 180, with a chance of being flipped. Due to the re-stitched prediction masks being binary, with white representing regions of interest and black representing areas not of interest, any rotation operation on the original ground truth mask tiles could result in inverting the areas of interest to be considered as regions not of interest by adding extra black space on the tile borders, thus introducing noise to the true performance. To overcome this, a function was deployed to trim the border regions of the augmented mask, thereby preserving the alignment between the training tile and the ground truth binary mask and reducing potential inconsistent tissue placement. For model training, the ViT-B/16 utilized a base learning rate of 5e-4 as the first hyperparameter. An AdamW optimizer was employed, along with a weight decay of 1e-4. Furthermore, a cosine learning rate was used, gradually adjusting the learning rate over time, which converges faster and potentially yields better results. Training used five epochs for warm-up, followed by a complete run of 50 epochs with a batch size of 64.
@@
Since the muscularis is the largest region within the colon, the tiles that were accepted correspond to random tiles; therefore, the tiles chosen for the model’s training purposes weren’t necessarily selected based on whether they contained plexus regions. After inputting each tile into the ViT model, the output consists of a pixel-wise raw logit, which is then converted to a probability between 0 and 1 after passing through a SoftMax activation function, representing the probability of the muscularis propria layer. The center region of the prediction tile is then retained and later used for re-stitching to form the full WSI prediction map. To evaluate the model’s predictive performance for each layer, a manual sweep operation was performed to ensure proper alignment and no missing regions between the ground truth segmentation binary map and the prediction segmentation binary mask. In terms of determining the most optimal threshold to accept a tile corresponding to a muscularis propria pixel, a threshold sweep of 0.01-0.99 was applied, and the threshold which produced the highest PIR and Dice coefficient was utilized. The effect of varying the prediction threshold on both Dice and PIR is summarized in Figure 5.

### v2

@@ Abstract
Hirschsprung Disease is characterized by the absence of ganglion cells in the myenteric plexus. Therefore, the correct identification of ganglion cells is crucial for diagnosing Hirschsprung disease. We introduce a three-stage analysis framework that mimics the pathologist’s diagnostic approach. The framework, based on a Vision Transformer model (ViT-B/16), sequentially segments the muscularis propria, segments the myenteric plexus, and detects ganglion cells within anatomically valid regions. 30 whole-slide images of colon tissue were used, each containing manual annotations of muscularis, plexus, and ganglion cells. A 5-fold cross-validation scheme was applied to each stage, along with resolution-specific tiling strategies and tailored postprocessing to ensure anatomical consistency. The proposed method achieved a Dice coefficient of 89.9% and a Plexus Inclusion Rate of 100% for muscularis segmentation. Plexus segmentation reached a recall of 94.8%, a precision of 84.2% and a Ganglia Inclusion Rate of 99.7%. For ganglion cells annotated with high certainty, the model achieved 62.1% precision and 89.1% recall. When considering all annotated ganglion cells, regardless of certainty level, the overall precision was 67.0%. These results indicate that ViT-based models are effective at leveraging global tissue context and capturing cellular morphology at small scales, even within complex histological tissue structures. This multi-stage methodology has great potential to support digital pathology workflows by reducing inter-observer variability and assisting in the evaluation of Hirschsprung disease. The clinical impact will be evaluated in future work with larger multi-center datasets and additional expert annotations.
@@
Hirschsprung disease (HD), a congenital birth defect defined by the absence of ganglionic cells within the myenteric plexus of the intestinal tract [7], affects 1/5000 neonatal patients worldwide [1]. Absence of ganglionic cells is due to defective migration, proliferation and differentiation of neural crest cells in the womb. It leads to malformation of enteric nerve cells that cause obstruction of the bowel. The patient presents with symptoms of severe constipation and features of intestinal obstruction [2]. Thus, it is necessary to identify the ganglion cells of the colon to confirm HD.
@@
After the onset of symptoms, traditional methods for diagnosing HD include a contrast enema radiograph, in which a barium contrast is inserted into the child’s rectum and x-ray images of the colon are taken to visualize the structure of different colon segments, thereby confirming the presence of malformed intestines. A second common method for diagnosing HD is a rectal biopsy, in which a section of tissue is extracted from the rectum and examined on a Whole Slide Image (WSI) by a trained pathologist. This method involves quantitatively assessing the presence of ganglion cells. Although both approaches are considered the gold standard for confirming HD, they have significant limitations, including sampling error, inter- and intra-rater variability among pathologists [8, 10], and substantial time requirements. More critically, rectal biopsies are used to determine the most distal healthy segment of the intestine that can be preserved during a surgical pull-through procedure, which is a primary treatment option. Misidentifying a healthy region as aganglionic may lead to unnecessary resection of functional bowel, while misclassifying an aganglionic segment as healthy can result in surgical failure, which is particularly consequential given the limited physiological reserve of newborn patients, highlighting the importance of automating the identification of ganglionic cells with high accuracy.
@@
Recent advances in computational imaging and digital pathology have enabled Artificial Intelligence (AI) methods to automate the classification [31], detection [32], and segmentation [33] of histopathology images across a wide range of organ systems. For instance, several Deep Learning (DL) algorithms have been utilized for computer-aided diagnosis, attempting to classify breast cancer lesions using WSIs [3] and human aortic and cardiac regions [5]. Other algorithms [4] have been deployed to segment images of bronchoscopes, further identifying lung-related diseases. Furthermore, object segmentation has also been widely applied to various medical images and diseases. Specifically, classifying the potential onset of autism spectrum disorder (ASD), including additional indicators of ASD and irregular brain structure [24, 25, 30]. Leveraging DL based computer vision for identifying ganglion cells is a promising direction. An effective strategy is to incorporate anatomical context into the workflow, since ganglion cells reside within the myenteric plexus, which itself is embedded in the muscularis layer. This motivates a multistage approach in which the muscularis is first identified, the plexus is then localized within that region, and ganglion cells are subsequently detected within the predicted plexus to confirm the presence of HD.
@@
Recent attempts [11, 12, 13, 14, 15] have been made to segment the muscularis layer, which encloses the myenteric plexus, first utilizing a shallow machine learning base model, K-means clustering [14]. This was then followed by deploying a Convolutional Neural Network (CNN), which produced a far greater accuracy and plexus inclusion rate when segmenting the muscularis layer [13, 14]. A colour-thresholding algorithm was then utilized for segmenting the myenteric plexus region and effectively detecting Carletinin-positive ganglia inside plexus regions. Finally, ganglia were segmented from each plexus region using intensity thresholding on the previously segmented Calretinin signal. This was followed by deploying a Linear Discriminant Analysis (LDA) model to identify true-ganglia from false-ganglia [14]. Overall, the results suggest strong discriminative ability, offering a promising avenue for automated DL. Other attempts have been made [6] that leverage the U-Net, a CNN algorithm, to similarly detect the presence of ganglion cells in the myenteric plexus, using WSIs from biopsy tissues. M. Duci et. al., [6] demonstrated innate capabilities of the CNN architecture, especially given its ability to extract low-level and high-level features throughout different layers, in addition to capturing local relationships through convolutions. Overall, the algorithm is able to learn rich, essential semantic features from the colon layers in the WSI dataset. Although U-Net is appropriate for learning pixel-wise local representations, it cannot learn long-range, global contextual relationships across different regions of the image, which could be critical for detecting ganglion cells dispersed throughout the WSI.
@@
Transformers were first introduced in Natural Language Processing (NLP), where their attention-based architecture demonstrated a strong ability to model long-range dependencies. Their success led to rapid adoption in computer vision, where Vision Transformers (ViTs) [17] have shown competitive and often superior performance to CNNs across general imaging benchmarks [17, 35, 36] and medical image analysis tasks [9, 15, 16, 30]. This is possible because ViTs have inherent self-attention [34], which enables them to model long-range dependencies and capture global contextual relationships across all parts of an image. Thereby, recognizing patterns that span large areas, such as tissue structure in WSI histopathology.
In this study, we introduce a three-stage semantic segmentation approach that mirrors the colon’s anatomical hierarchy. The first stage identifies the muscularis layer, which provides the outer boundary within which the myenteric plexus regions are located. The second stage segments the myenteric plexus regions inside the predicted muscularis. The third stage then detects ganglion cells "objects" within the plexus region, enabling assessment of ganglion cell presence for HD diagnosis. The study evaluates a three-stage ViT-based pipeline that processes WSIs in a tile-based manner. In each stage, the WSI resolution is adjusted so that a 224224 tile captures an appropriate field of view (FOV) for the anatomical structure of interest while still preserving the level of detail needed for accurate segmentation. This allows the model to access broad contextual information for larger structures, such as the muscularis, and finer morphological features for smaller structures, such as ganglion cells.
Performance is assessed at each stage of the pipeline and for the combined workflow, and results are compared with previously published methods that used the same dataset [11, 13, 14]. This framing highlights the contribution of each stage to the overall identification task and enables a direct comparison with earlier approaches that relied on shallow learning or CNN-based models.
@@ 2.1 Dataset
The dataset used in this research study originates from the Children’s Hospital of Eastern Ontario (CHEO), comprising 30 WSIs corresponding to colon sections from 26 different patients, all of whom were diagnosed with HD. All high-resolution images were extracted from prepared tissue slides after being scanned using the Aperio ScanScope CS (Aperio Technologies) at a 20 magnification level (0.50 /pixel). All 30 WSIs were saved in an SVS format. For each layer, there exists ground truth annotations that a graduate researcher manually segmented (and approved by a pediatric pathologist) [14] (Fig. 1). For instance, the muscularis propria was manually delineated to represent the ground-truth segmentation (Fig. 2A), the myenteric plexus region was roughly manually segmented (a visually noticeable amount of tissue around the plexus regions was also included) (Fig. 2B). The ganglion cells were also roughly segmented with a confidence level accompanies each ganglion cell annotation. A high-confidence level indicates strong certainty that the annotated object is a ganglion cell (Fig. 2C). In contrast, a low-confidence level reflects that the object is believed to be a ganglion cell but with some uncertainty (Fig. 2D), due to the difficulty in recognizing and selecting physiological features of a ganglion cell [15].
@@ 2.2.1 Muscularis Propria
Preprocessing for the muscularis propria region followed a three-step procedure. First, WSIs were downsampled from the original 20 (0.50 /pixel) resolution to 5 /pixel (equivalent to a 10 reduction in spatial resolution). At this lower resolution, each 224224 tile covers a larger physical FOV, which is appropriate for capturing the broad, continuous architecture of the muscularis propria. Second, colour normalization was applied using the Macenko method [18] to mitigate staining variation across slides. Third, each WSI was tiled into 224224-pixel subimages, from which 1000 tiles were randomly selected per slide (total of 30,000 tiles). These tiles enabled the model to learn characteristic muscularis patterns across the dataset. An example of the corresponding ground truth annotations is shown in Fig. 1B.
@@ 2.2.2 Myenteric Plexus
Preprocessing for the myenteric plexus followed the same general workflow but used a resolution tailored to plexus morphology. WSIs were downsampled from the original 20 (0.50 /pixel) resolution to 2.5 /pixel (equivalent to a 5 reduction in spatial resolution). At this level, each 224224 tile spans a FOV wide enough to capture elongated plexus structures while still preserving sufficient detail to distinguish them from surrounding muscularis tissue. Macenko normalization [18] was applied for stain harmonization. WSIs were then tiled into 224224 pixel patches, and tiles were retained only if they contained at least one pixel from the muscularis propria region. Since the myenteric plexus is anatomically confined to the muscularis, this ensured that only relevant tissue regions were included. Examples of plexus annotations are shown in Fig. 1C and Fig. 3B.
@@ 2.2.3 Ganglion Cells
Preprocessing for ganglion cell analysis also followed this framework, but used a higher-resolution image appropriate for cellular-level features. WSIs were downsampled from the original 20 (0.50 /pixel) resolution to 1.0 /pixel (equivalent to a 2 reduction in spatial resolution). This higher resolution allows each 224224 tile to preserve fine morphological details necessary for identifying ganglion cells. Macenko normalization [18] was again applied. WSIs were tiled into 224224 pixel patches, and tiles were retained only if they contained at least one pixel from the myenteric plexus regions. Because ganglion cells occur exclusively within the plexus, this ensured that model training was restricted to anatomically relevant regions. Examples of ganglion cells are shown in Fig. 2C and Fig. 2D.
@@ 2.3.1 Model Architecture
The three-stage analysis framework was based on a pretrained ViT-B/16 backbone [17]. Specifically, each 224224 tile was divided into fixed-sized patches (1616) that were linearly projected to patch embeddings and then fed to the transformer encoder. The encoder at each stage produced a global contextual representation, which was used by a task-specific decoder to generate the corresponding prediction map. Because the three stages operate independently, separate ViT-based models were fine-tuned for muscularis segmentation, plexus segmentation, and ganglion cell detection. At each stage, the decoder received the encoded representations and produced a dense, pixel-wise output for that anatomical layer. An overview of the complete three-stage pipeline is provided in Fig. 4.
@@ 2.3.2 Muscularis Propria
To ensure robust segmentation of the muscularis propria and to reduce the risk of overfitting given the limited number of WSIs, a 5-fold cross-validation strategy was employed at the WSI level. The dataset was partitioned into five disjoint folds, each containing 6 WSIs. In each iteration, four folds were used for model training, while the remaining fold was held out for evaluation. From each WSI in the training folds, 1000 tiles were randomly extracted and used for model optimization, resulting in 24,000 training tiles per fold. Tiles extracted from the held-out WSIs were not used to update model parameters and were employed only to monitor validation performance during training and to guide early stopping.
Final model evaluation was performed at the WSI level rather than at the tile level. For evaluation, segmentation was carried out using overlapping tiles across each held-out WSI, with a tile size of pixels and a stride of 112 pixels. Each tile was segmented by the ViT model, and only the central pixel region of each predicted tile was retained. These central predictions were then stitched together to reconstruct a full-resolution WSI segmentation map. Quantitative performance metrics were computed by comparing the reconstructed WSI-level prediction maps with the corresponding ground truth segmentation masks. This process was repeated across folds so that each WSI served as the evaluation case exactly once. The evaluation process was used in the other two stages as well (myenteric plexus segmentation and ganglion cells detection).
@@
To increase dataset diversity, the tile training sets were augmented using four methods: scaling, random rotation, horizontal flipping, and vertical flipping. The scaling operation simulated zoom effects by randomly shrinking the visible region by at least 50% before resizing the tile back to 224224 pixels. This augmentation enabled the model to learn from tiles representing different effective magnifications. Rotation entailed rotating the tiles by 90; horizontal and vertical flipping consisted of flipping the images by 180, with a chance of being flipped. Due to the re-stitched prediction masks being binary, with white representing regions of interest and black representing areas not of interest, any rotation operation on the original ground truth mask tiles could result in inverting the areas of interest to be considered as regions not of interest by adding extra black space on the tile borders, thus introducing noise to the true performance. To overcome this, a function was deployed to trim the border regions of the augmented mask, thereby preserving the alignment between the training tile and the ground truth binary mask and reducing potential inconsistent tissue placement.
For model training, the ViT-B/16 utilized a base learning rate of 5e-4 as the first hyperparameter. An AdamW optimizer was employed, along with a weight decay of 1e-4. Furthermore, a cosine learning rate was used, gradually adjusting the learning rate over time, which converges faster and potentially yields better results. Training used five epochs for warm-up, followed by a complete run of 50 epochs with a batch size of 64.
@@
Since the muscularis is the largest region within the colon, the tiles that were accepted correspond to random tiles; therefore, the tiles chosen for the model’s training purposes weren’t necessarily selected based on whether they contained plexus regions. After inputting each tile into the ViT model, the output consists of a pixel-wise raw logit, which is then converted to a probability between 0 and 1 after passing through a SoftMax activation function, representing the probability of the muscularis propria layer. The center region of the prediction tile is then retained and later used for re-stitching to form the full WSI prediction map. To evaluate the model’s predictive performance for each layer, a manual sweep operation was performed to ensure proper alignment and no missing regions between the ground truth segmentation binary map and the prediction segmentation binary mask.